An electron is in the n = 3 state. What is the smallest possible value for the angle between the z component of the orbital angular momentum and the orbital angular momentum?
Answer #1

The smallest value for \theta corresponds to the largest value of cos\theta.

Using

cos\theta =\sqrt{\frac{l}{l+1}}

For n = 3 largest allowed value of l = 2

hence

cos\theta =\sqrt{\frac{2}{2+1}}=0.8165

\theta =35.26^{o}